# Optimal Denial-of-Service Attacks Against Partially-Observable Real-Time Monitoring Systems

### v1

@@ IV-A Single-Source Scenario
Fig. 2 illustrates the evolution of the average reward under the optimal jamming policy derived in Theorem 2 and under the uniform random jamming policy as a function of $\lambda$. We supposed that $\lambda$ varies from 00 to $10$ with step size $0.001$, and we considered the following parameters $p=q=0.9$ and $r=0.1$. We can observe that the average reward under the optimal solution decreases as $\lambda$ grows and converges to a fixed value. Indeed, when $\lambda$ increases, the cost of jamming becomes more substantial in terms of energy, which reduces the performance of the optimal solution. We also observe that the uniform random jamming policy is sub-optimal, which corroborates our theoretical results.
@@ V Conclusion
In this paper, we studied the problem of denial-of-service attacks against status updating in a partially observable environment. The target system was modeled by Markov chains and i.i.d. channels, and the performance of status updating was measured based on the AoII. We derived the structures of the optimal jamming policies in single and multi-source scenarios. Specifically, for the single-source scenario, we proved that the optimal jamming policy is a threshold policy, and developed a low-complexity algorithm to find the optimal threshold value. Additionally, for the multi-source scenario, we adopted the Lagrangian relaxation approach to develop a low-complexity and well-performing policy, dubbed Whittle index jamming policy. Finally, our numerical results showed that the proposed policy for both scenarios can outperform random policies, corroborating thus our theoretical findings.

### v2

@@ IV-A Single-Source Scenario
Fig. 2 illustrates the evolution of the average reward under the optimal jamming policy derived in Theorem 2 and under the uniform random jamming policy as a function of $\lambda$. We supposed that $\lambda$ varies from 00 to $10$ with step size $0.001$, and we considered the following parameters $p=q=0.9$ and $r=0.1$. As $\lambda$ increases, the average reward for the optimal solution declines and eventually stabilizes at a constant value. This is because higher values of $\lambda$ result in a greater energy cost for jamming, which in turn diminishes the performance of the optimal strategy.. We also observe that the uniform random jamming policy is sub-optimal, which corroborates our theoretical results.
@@ V Conclusion
In this paper, we studied the problem of denial-of-service attacks against status updating in a partially observable environment. The target system was modeled by Markov chains and i.i.d. channels, and the performance of status updating was measured based on the AoII. We derived the structures of the optimal jamming policies in single and multi-source scenarios. For the single-source case, we demonstrated that the optimal jamming strategy follows a threshold policy and designed a low-complexity algorithm to determine the optimal threshold. Additionally, for the multi-source scenario, we adopted the Lagrangian relaxation approach to develop a low-complexity and well-performing policy, dubbed Whittle index jamming policy. Finally, our numerical results showed that the proposed policy for both scenarios can outperform random policies, corroborating thus our theoretical findings.